# CORE4D: A 4D Human-Object-Human Interaction Dataset for Collaborative Object REarrangement

### v1

@@ Abstract
Understanding how humans cooperatively rearrange household objects is critical for VR/AR and human-robot interaction. However, in-depth studies on modeling these behaviors are under-researched due to the lack of relevant datasets. We fill this gap by presenting CORE4D, a novel large-scale 4D human-object-human interaction dataset focusing on collaborative object rearrangement, which encompasses diverse compositions of various object geometries, collaboration modes, and 3D scenes. With 1K human-object-human motion sequences captured in the real world, we enrich CORE4D by contributing an iterative collaboration retargeting strategy to augment motions to a variety of novel objects. Leveraging this approach, CORE4D comprises a total of 11K collaboration sequences spanning 3K real and virtual object shapes. Benefiting from extensive motion patterns provided by CORE4D, we benchmark two tasks aiming at generating human-object interaction: human-object motion forecasting and interaction synthesis. Extensive experiments demonstrate the effectiveness of our collaboration retargeting strategy and indicate that CORE4D has posed new challenges to existing human-object interaction generation methodologies. Our dataset and code are available at https://github.com/leolyliu/CORE4D-Instructions.
@@ 1 Introduction
Humans frequently rearrange household items through multi-person collaboration , such as moving a table or picking up an overturned chair together. Analyzing and synthesizing these diverse collaborative behaviors could be widely applicable in VR/AR, human-robot interaction [77, 55, 54], and dexterous [10, 90, 76, 99] and humanoid [53, 18, 86, 42] manipulation. However, understanding and modeling these interactive motions have been under-researched due to the lack of large-scale, richly annotated datasets. Most existing human-object and hand-object interaction datasets focus on individual behaviors [71, 3, 49, 94, 20, 34, 100, 40, 46, 98] and two-person handovers [96, 81, 48]. But these datasets typically encompass a limited number of object instances, thus struggling to support generalizable interaction understanding across diverse object geometries. Scaling up precise human-object interaction data is challenging. While vision-based human-object motion tracking methods [3, 85] have made significant progress, they still struggle with low fidelity in severe occlusion, common in multi-human collaboration scenes. However, mocap [34, 40] is expensive and hard to scale up to cover a large number of objects to be rearranged. We want to curate a large-scale category-level human-object-human (HOH) interaction dataset with high motion quality in a cost-efficient manner.
@@
We observe that HOH collaborations mainly vary in two aspects including the temporal collaboration patterns of two humans and the spatial relations between human and object. The temporal collaboration patterns could vary in many ways depending on scene complexity, motion range, and collaboration mode. In contrast, the spatial relations between human and object tend to possess strong homogeneity when facing objects from the same category, e.g., two persons holding the two sides of a chair. This allows retargeting interactions involving one specific instance to another using automatic algorithms, avoiding the need to capture interactions with thousands of same-category objects in the real world. The above observations make it possible for us to leverage expensive motion capture systems to capture only humans’ diverse temporal collaboration patterns while leaving the richness of human-object spatial relations to automatic spatial retargeting algorithms.
@@
Using these insights, we build a large-scale dataset, CORE4D, encompassing a wide range of human-object interactions for collaborative object rearrangement. CORE4D includes various types of household objects, collaboration modes, and 3D environments. Our data acquisition strategy combines mocap-based capturing and synthetic retargeting, allowing us to scale the dataset effectively. The retargeting algorithm transfers spatial relation between human and object to novel object geometries while preserving temporal pattern of human collaboration. As a result, CORE4D includes 1K real-world motion sequences (CORE4D-Real) paired with videos and 3D scenes, and 10K synthetic collaboration sequences (CORE4D-Synthetic) covering 3K diverse object geometries.
@@
We benchmark two tasks for generating human-object collaboration: (1) motion forecasting [12, 88] and (2) interaction synthesis [73, 40] on CORE4D, revealing challenges in modeling human behaviors, enhancing motion naturalness, and adapting to new object geometries. Ablation studies demonstrate the effectiveness of our hybrid data acquisition strategy, and the quality and value of CORE4D-Synthetic, highlighting its role in helping to improve existing motion generation methods.
@@
In summary, our main contributions are threefold: (1) We present CORE4D, a large-scale 4D HOH interaction dataset for collaborative object rearrangement. (2) We propose a novel hybrid data acquisition methodology, incorporating real-world data capture and synthetically collaboration retargeting. (3) We benchmark two tasks for collaboration generation, revealing new challenges and research opportunities.
@@ 2.1 Human-object Interaction Datasets
Tremendous progress has been made in constructing human-object interaction datasets. To study how humans interact with 3D scenes, various widely-used datasets record human movements and surrounding scenes separately, regarding objects as static [4, 66, 28, 80, 27, 44, 17, 102, 104, 26, 79, 2, 31, 91, 16, 25, 72, 92, 35, 97] or partially deformable [43] without pose changes. For dynamic objects, recent works [5, 71, 15, 3, 32, 34, 100, 40, 81, 48, 105, 101, 45] have captured human-object interaction behaviors with different focuses. Table 1 generally summarizes the characteristics of 4D human-object-interaction datasets. To support research for vision-based human-object motion tracking and shape reconstruction, a line of datasets [15, 3, 32, 34, 100, 105, 101, 45] presents human-object mesh annotations with multi-view RGB or RGBD signals. With the rapid development of human-robot cooperation, several works [5, 71, 81, 48] focus on specific action types such as grasping [71] and human-human handover [5, 81, 48]. Our dataset uniquely captures multi-person and object collaborative motions, category-level interactions, and both egocentric and allocentric views, offering comprehensive features with the inclusion of both real and synthetic datasets.
@@ 2.2 Human Interaction Retargeting
Human interaction retargeting focuses on how to apply human interactive motions to novel objects in human-object interaction scenarios. Existing methodologies [37, 65, 94, 67, 83, 8, 33] are object-centric, which propose first finding contact correspondences between the source and the target objects and then adjusting human motion to touch specific regions on the target object via optimization. As crucial guidance of the result, contact correspondences are discovered by aligning either surface regions [65, 94, 83], spatial maps [37, 33], distance fields [8], or neural descriptor fields [67] between the source and the target objects, which are all limited to objects with similar topology and scales. Our synthetic data generation strategy incorporates object-centric design [94] with novel human-centric contact selection, creating a chance to adapt to these challenging objects using human priors.
@@ 2.3 Human-object Interaction Generation
Human-object interaction generation is an emerging research topic that aims to synthesize realistic human-object motions conditioned on surrounding 3D scenes, known object trajectories, or action types. To generate humans interacting with static 3D scenes, POSA [29] and COINS [106] synthesize static human poses with CVAE [68], while a line of work [69, 22, 84, 70, 104, 38, 103] further presents dynamic human motions by auto-regressive manners [69, 104], diffusion models [38], or two-stage designs that first generates start and end poses and then interpolates motion in-between [22, 84, 70, 103]. InterDiff [88] and OMOMO [40] further fulfill this task for dynamic objects. To generate human-object interaction under specific action descriptions, recent works [19, 39, 63, 82, 89] extract text features with pretrained CLIP [64] encoders or LLMs [57, 74] and use them to guide diffusion models [30].
@@ 3 Constructing CORE4D
CORE4D is a large-scale 4D human-object-human interaction dataset acquired in a novel hybrid scheme, comprising CORE4D-Real and CORE4D-Synthetic. CORE4D-Real is captured (Section 3.1) and annotated (Section 3.2) from authentic collaborative scenarios. It provides human-object-human poses, allocentric RGB-D videos, egocentric RGB videos, and 2D segmentation across 1.0K sequences accompanied by 37 object models. To augment spacial relation between human and object, we present an innovative collaboration retargeting technique in Section 3.3, integrating CORE4D-Real with CORE4D-Synthetic, thereby expanding our collection with an additional 10K sequences and 3K rigid objects. Detailed characteristics such as data diversities are demonstrated in Section 3.4.
@@ 3.1 CORE4D-Real Data Capture
Figure 1: Data capturing system. (a) demonstrates the wearing of mocap suits and the positioning of the egocentric camera. (b) shows an object with four markers. (c) illustrates the data capturing system and camera views.
Figure 2: Dataset overview.
To collect precise human-object motions with visual signals, we set up a hybrid data capturing system shown in Fig. 1, consisting of an inertial-optical mocap system, four allocentric RGB-D cameras and a camera worn by persons for egocentric sensing. The frequency of our system is 15 FPS.
@@
Visual Sensors. Kinect Azure DK cameras are integrated to capture allocentric RGB-D signals, and an Osmo Action3 is utilized to capture egocentric color videos. The resolution of all the visual signals is 1920x1080. Cameras are calibrated by the mocap system and synchronized via timestamp. Details on camera calibration and synchronization are provided in the supplementary material.
@@
Privacy Protection. To ensure participant anonymity, blurring is applied to faces [58] in RGB videos, and fake facial meshes are generated via SMPL-X [61]. The participants all consented to releasing CORE4D, and were also notified of their right to have their data removed from CORE4D at any time.
@@ 3.2 CORE4D-Real Data Annotation
Human Mesh Acquisition. Aligning with existing dataset efforts [34, 40], we retarget the BVH [52] human skeleton to the widely-used SMPL-X [61]. SMPL-X [61] formulates a human mesh as $D_{\text{smplx}}=M(\beta,\theta)$. The body shape $\beta\in\mathbb{R}^{10}$ are optimized to fit the constraints on manually measured human skeleton lengths. With $\beta$ computed, we optimize the full-body pose $\theta\in\mathbb{R}^{159}$ with the loss function:
@@
where $\mathcal{L}_{\text{reg}}$ ensures the simplicity of the results and prevents unnatural, significant twisting of the joints. $\mathcal{L}_{j\text{3D}}$ and $\mathcal{L}_{j\text{Ori}}$ encourage the rotation of joints and the global 3D positions to closely match the ground truth. $\mathcal{L}_{h\text{3D}}$ and $\mathcal{L}_{h\text{Ori}}$ guide the positioning and orientation of the fingers. $\mathcal{L}_{\text{smooth}}$ promotes temporal smoothness. $\mathcal{L}_{\text{contact}}$ encourages realistic contact between the hands and objects. Then using SMPL-X [61] $M(\beta,\theta,\Phi):\mathbb{R}^{|\theta|\times|\beta|}\mapsto\mathbb{R}^{3N}$ to generate human mesh. Details on the loss functions are presented in the supplementary material.
2D Mask Annotation. We offer automatic 2D segmentation for individuals and the manipulated object to aid in predictive tasks like vision-based human-object pose estimation [3, 85]. We first use DEVA [9] to segment human and object instances in a captured interaction image with text prompts. Then, we render human and object meshes separately on each image and select the instance with the highest Intersection-over-Union (IoU) for mask annotation.
Figure 3: Collaboration retargeting pipeline. We propose a collaboration retargeting algorithm by iteratively refining interaction motion. The algorithm takes input source-target pair. First, we sample contact candidates from whole CORE4D-Real contact knowledge on source. For each contact candidate, we apply contact retargeting to propagate contact candidates to contact constraints on target. Sampled motion from CORE4D-Real provides a high-level collaboration pattern, together with low-level contact constraints, we obtain interaction candidates from interaction retargeting. Then, the human pose discriminator selects the optimal candidates, prompting a contact constraints update via beam search. After multiple iterations, the process yields augmented interactions. This iterative mechanism can effectively get a reasonable one from numerous contact constraints and ensures a refined interaction, enhancing the dataset’s applicability across various scenarios.
@@ 3.3 CORE4D-Synthetic Data Generation
where fidelity loss $\mathcal{L}_{\text{j}}$ evaluates the difference in human joint positions before and after the transfer, contact loss $\mathcal{L}_{c}$ computes the difference between human-object contact regions and the contact constraints, $\mathcal{L}_{\text{spat}}$ and $\mathcal{L}_{\text{smooth}}$ ensures the smoothness of human motion. Details on the loss designs and their motivations are provided in the supplementary material.
Human-centric Contact Selection. Selecting reasonable contact constraints efficiently is challenging due to their large scale and the time-consuming interaction retargeting. We address this challenge by developing a beam search algorithm to select contact constraints from a human-centric perspective. Specifically, we train a human pose discriminator inspired by GAN-based motion generation strategies [93, 87]. To train the discriminator, we build a pairwise training dataset, with each pair consisting of one positive human pose sample and one negative sample. Positive samples are encouraged to get higher scores than negative ones. We use CORE4D-Real as positive samples. We add 6D pose noise $\Delta(\alpha,\beta,\gamma,x,y,z)$ on target motion, and regard corresponding human motions generated by contact-guided interaction retargeting as negative samples. The loss function is:

### v2

@@ Abstract
Understanding how humans cooperatively rearrange household objects is critical for VR/AR and human-robot interaction. However, in-depth studies on modeling these behaviors are under-researched due to the lack of relevant datasets. We fill this gap by presenting CORE4D, a novel large-scale 4D human-object-human interaction dataset focusing on collaborative object rearrangement, which encompasses diverse compositions of various object geometries, collaboration modes, and 3D scenes. With 1K human-object-human motion sequences captured in the real world, we enrich CORE4D by contributing an iterative collaboration retargeting strategy to augment motions to a variety of novel objects. Leveraging this approach, CORE4D comprises a total of 11K collaboration sequences spanning 3K real and virtual object shapes. Benefiting from extensive motion patterns provided by CORE4D, we benchmark two tasks aiming at generating human-object interaction: human-object motion forecasting and interaction synthesis. Extensive experiments demonstrate the effectiveness of our collaboration retargeting strategy and indicate that CORE4D has posed new challenges to existing human-object interaction generation methodologies.
CORE4D is a large-scale diverse human-object-human interaction dataset for collaborative object rearrangement, encompassing real-world and synthetic branches. CORE4D-Real captures 1K human-object-human mesh sequences with allocentric and egocentric visual signals, while CORE4D-Synthetic retargets real-world data onto 3K virtual object shapes formulating 10K motion sequences.
11footnotetext: Equal contribution.22footnotetext: Corresponding author.
@@ 1 Introduction
Humans frequently rearrange household items through multi-person collaboration , such as moving a table or picking up an overturned chair together. Analyzing and synthesizing these diverse collaborative behaviors could be widely applicable in VR/AR, human-robot interaction [103, 73, 72], dexterous manipulation [11, 121, 104, 134] and humanoid manipulation [71, 20, 117, 50]. However, understanding and modeling these interactive motions have been limited due to the lack of large-scale, richly annotated datasets. Most existing human-object and hand-object interaction datasets focus on individual behaviors [92, 3, 57, 128, 24, 41, 135, 53, 59, 132] and two-person handovers [129, 109, 58]. But these datasets typically encompass a limited number of object instances, thus struggling to support generalizable interaction understanding across diverse object shapes. Scaling up precise human-object interaction data is challenging. While vision-based human-object motion tracking methods [114, 141, 115, 116] have advanced, they still struggle with low fidelity due to severe occlusions, which are common in multi-human collaborations. Additionally, mocap [41, 53] is expensive and hard to scale up to cover numerous objects involved in rearrangement. Our goal is to curate a large-scale category-level human-object-human (HOH) interaction dataset with high quality in a cost-efficient manner.
@@
We observe that HOH collaborations mainly vary in two aspects: the temporal collaboration patterns of two humans and the spatial relations between human and object. The temporal collaboration patterns could vary widely depending on scene complexity, motion range, and collaboration mode. In contrast, the spatial relations between human and object tend to possess strong homogeneity when facing objects from the same category, e.g., two persons holding opposite sides of a chair. This allows for retargeting interactions involving one specific instance to another using automatic algorithms, avoiding the need to capture interactions with thousands of same-category objects in the real world. The above observations enable us to leverage expensive motion capture systems to capture only humans’ diverse temporal collaboration patterns, while relying on automatic spatial retargeting algorithms to enrich human-object spatial relations.
@@
Using these insights, we build a novel large-scale dataset, CORE4D, encompassing a wide range of human-object interactions for collaborative object rearrangement. CORE4D includes various types of household objects, collaboration modes, and 3D environments. Our data acquisition strategy combines mocap-based capturing and synthetic retargeting, allowing us to scale the dataset effectively. The retargeting algorithm transfers spatial relation between human and object to novel object geometries while preserving temporal pattern of human collaboration. As a result, CORE4D includes 1K real-world motion sequences (CORE4D-Real) paired with videos and 3D scenes, as well as 10K synthetic collaboration sequences (CORE4D-Synthetic) covering 3K diverse object shapes.
@@
We benchmark two tasks for generating human-object collaboration: (1) motion forecasting [14, 119] and (2) interaction synthesis [97, 53] on CORE4D, revealing challenges in modeling human behaviors, enhancing motion naturalness, and adapting to new object geometries. Ablation studies demonstrate the effectiveness of our hybrid data acquisition strategy, and the quality and value of CORE4D-Synthetic, highlighting its role in helping to improve existing motion generation methods. We further retarget interactions in CORE4D onto Unitree H1 [100] humanoid robot and use them to train humanoid box-lifting policies, showcasing the values of CORE4D in robot interaction skill learning.
@@
In summary, our main contributions are threefold: (1) We present CORE4D, a large-scale 4D HOH interaction dataset for collaborative object rearrangement. (2) We propose a novel hybrid data acquisition method, incorporating real-world data capture and synthetic collaboration retargeting. (3) We benchmark two tasks for collaboration generation, revealing new challenges and research opportunities.
@@ 2.1 Human-object Interaction Datasets
Tremendous progress has been made in constructing human-object interaction datasets. To study how humans interact with 3D scenes, various widely-used datasets record human movements and surrounding scenes separately, treating objects as static [4, 86, 32, 106, 33, 55, 18, 137, 139, 30, 107, 2, 38, 123, 19, 31, 95, 124, 43, 130, 42] or partially deformable [54] without pose changes. For dynamic objects, recent works [5, 13, 6, 45, 48, 98, 92, 17, 3, 39, 41, 135, 53, 109, 58, 140, 136, 56, 141, 63, 126] have captured human-object interaction behaviors with varying focuses. Table 1 generally summarizes the characteristics of 4D human-object-interaction datasets. To support research for vision-based human-object motion tracking and shape reconstruction, a line of datasets [17, 3, 39, 41, 135, 140, 136, 56] present human-object mesh annotations with multi-view RGB or RGBD signals. With the rapid development of human-robot cooperation, several works [5, 92, 109, 58] focus on specific action types such as grasping [92] and human-human handover [5, 109, 58]. Our dataset uniquely captures multi-person and object collaborative motions, category-level interactions, and both egocentric and allocentric views, offering comprehensive features with the inclusion of both real and synthetic datasets.
@@ 2.2 Human Interaction Retargeting
Human interaction retargeting focuses on applying human interactive motions to novel objects in human-object interaction scenarios. Existing methodologies [46, 85, 128, 88, 111, 9, 40, 115] are object-centric. They propose first finding contact correspondences between the source and the target objects and then adjusting human motion to touch specific regions on the target object via optimization. As crucial guidance for the result, contact correspondences are discovered by aligning either surface regions [85, 128, 111, 115], spatial maps [46, 40], distance fields [9], or neural descriptor fields [88] between the source and the target objects. These methods are all limited to objects with similar topology and scales. Our synthetic data generation strategy incorporates object-centric design [128] with novel human-centric contact selection, allowing adaptation to challenging objects using human priors.
@@ 2.3 Human-object Interaction Generation
Human-object interaction generation is an emerging research topic that aims to synthesize realistic human-object motions conditioned on surrounding 3D scenes, known object trajectories, or action types. To generate 3D human mesh snapshots interacting with scenes, POSA [34] and COINS [142] propose to leverage CVAE [89], while DreamHOI [145] provides an iterative NeRF [69] optimization approach. To further synthesize interactive motions, a line of works [91, 27, 112, 93, 139, 49, 138, 70] present auto-regressive manners [91, 139], diffusion models [49], or two-stage designs that first generates start and end poses and then interpolates motion in-between [27, 112, 93, 138]. Beyond static objects, a line of works further model object movements and generate integrated human-object interactions using diffusion models [119, 53] and GCN [122]. To generate human-object interactions under action descriptions, recent works [22, 52, 81, 110, 120, 113, 108, 131, 23, 42, 90] extract text features with pretrained CLIP encoders [83, 108, 131, 42, 90] or LLM planners [75, 99, 113, 23], using them to guide diffusion models [37].
@@ 3 Constructing CORE4D
CORE4D is a large-scale 4D human-object-human interaction dataset acquired in a novel hybrid scheme, comprising CORE4D-Real and CORE4D-Synthetic. CORE4D-Real is captured (Section 3.1) and annotated (Section 3.2) from authentic collaborative scenarios. As shown in Figure 2, it provides human-object-human poses, egocentric RGB videos, allocentric RGB-D videos, and 2D segmentations across 1.0K sequences accompanied by 37 object models. To augment spacial relation between human and object, we present an innovative collaboration retargeting technique in Section 3.3. This technique integrates CORE4D-Real with CORE4D-Synthetic, thereby expanding our collection with an additional 10K sequences and 3K rigid objects. Detailed characteristics such as data diversities are discussed in Section 3.4.
@@ 3.1 CORE4D-Real Data Capture
Figure 1: CORE4D-Real data capturing system. (a) demonstrates the wearing of mocap suits and the positioning of the egocentric camera. (b) shows an object with four markers. (c) illustrates the data capturing system and camera views.
Figure 2: CORE4D-Real data modality overview.
To collect precise human-object motions with visual signals, we set up a hybrid data capturing system shown in Fig. 1, consisting of an inertial-optical mocap system, four allocentric RGB-D cameras and a camera worn by participants for egocentric sensing. The system operates at 15 FPS.
@@
Visual Sensors. Kinect Azure DK cameras are integrated to capture allocentric RGB-D signals, and an Osmo Action3 is utilized to capture egocentric color videos. The resolution of all the visual signals is 1920x1080. Cameras are calibrated by the mocap system and synchronized via timestamp. Details on camera calibration and synchronization are in the appendix.
@@
Privacy Protection. To ensure participant anonymity, blurring is applied to faces [76] in RGB videos. The participants all consented to releasing CORE4D, and were notified of their right to remove their data from CORE4D at any time.
@@ 3.2 CORE4D-Real Data Annotation
Human Mesh Acquisition. Aligning with existing datasets [41, 53], we retarget BVH [67] human skeletons to SMPL-X [79]. SMPL-X [79] formulates a human mesh as $D_{\text{smplx}}=M(\beta,\theta)$. The body shape $\beta\in\mathbb{R}^{10}$ are optimized to fit the constraints on manually measured human skeleton lengths. With $\beta$ computed, we optimize the full-body pose $\theta\in\mathbb{R}^{159}$ with the loss function:
@@
where $\mathcal{L}_{\text{reg}}$ ensures the simplicity of the results and prevents unnatural, significant twisting of the joints. $\mathcal{L}_{j\text{3D}}$ and $\mathcal{L}_{j\text{Ori}}$ encourage the rotation of joints and the global 3D positions to closely match the ground truth. $\mathcal{L}_{h\text{3D}}$ and $\mathcal{L}_{h\text{Ori}}$ guide the positioning and orientation of the fingers. $\mathcal{L}_{\text{smooth}}$ promotes temporal smoothness. $\mathcal{L}_{\text{contact}}$ encourages realistic contact between the hands and objects. Then using SMPL-X [79] $M(\beta,\theta,\Phi):\mathbb{R}^{|\theta|\times|\beta|}\mapsto\mathbb{R}^{3N}$ to generate human mesh. Details on loss functions are in the appendix.
2D Mask Annotation. We offer automatic 2D segmentation for individuals and the manipulated objects to aid in predictive tasks like vision-based human-object pose estimation [3, 114]. We first use DEVA [10] to segment human and object instances in a captured interaction image with text prompts. Then, we render human and object meshes separately on each image and select the instance with the highest Intersection-over-Union (IoU) for mask annotation.
Figure 3: Collaboration retargeting pipeline. We propose a collaboration retargeting algorithm by iteratively refining interaction motion. The input is a source-target pair. First, we sample contact candidates from whole CORE4D-Real contact knowledge on source. For each contact candidate, we apply contact retargeting to propagate contact candidates to contact constraints on target. Sampled motion from CORE4D-Real provides a high-level collaboration pattern, together with low-level contact constraints, we obtain interaction candidates from interaction retargeting. Then, the human pose discriminator selects the optimal candidates, prompting a contact constraints update via beam search. After multiple iterations, the process yields augmented interactions. This iterative mechanism can effectively get a reasonable one from numerous contact constraints and ensures a refined interaction, enhancing the dataset’s applicability across various scenarios.
@@ 3.3 CORE4D-Synthetic Data Generation
where fidelity loss $\mathcal{L}_{\text{j}}$ evaluates the difference in human joint positions before and after the transfer, contact loss $\mathcal{L}_{c}$ computes the difference between human-object contact regions and the contact constraints, $\mathcal{L}_{\text{spat}}$ and $\mathcal{L}_{\text{smooth}}$ ensures the smoothness of human motion. Details on the loss designs are in the appendix.
Human-centric Contact Selection. Selecting reasonable contact constraints efficiently is challenging due to their large scales and the time-consuming interaction retargeting. We address this challenge by developing a beam search algorithm to select contact constraints from a human-centric perspective. Specifically, we train a human pose discriminator inspired by GAN-based motion generation works [125, 118]. To train it, we build a pairwise training dataset, with each pair consisting of one positive human pose sample and one negative one. Positive samples are encouraged to get higher scores than negative ones. We use CORE4D-Real as positive samples. We add 6D pose noise $\Delta(\alpha,\beta,\gamma,x,y,z)$ on target motion, and regard corresponding human motions generated by contact-guided interaction retargeting as negative samples. The loss function is: